OpenStudy (anonymous):

(xx)(xx)(xx)(xx) = xxxxxxxx is not true (xx)(xx)(xx)(xx) = xxxxxxxx is not true @Mathematics

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

xx times xx = xxxxx right? 5*5= 25 25*5=125*5=625!

OpenStudy (anonymous):

(x2)4 = (x2)(x2)(x2)(x2) = (xx)(xx)(xx)(xx) = xxxxxxxx = x8 with numbers 2squared =4 4*4=16 4to the 8th power =65536

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

please answer!